OpenStudy (anonymous):

The sum of the quotients of 25x25x and the quotient of 8x2x is

OpenStudy (anonymous):

\[25x+25x ?\]

OpenStudy (anonymous):

if this the case \[25x+25x=x(25+25)=50x\]

OpenStudy (anonymous):